Hey, My math question is
"A rectangulat garden has an area of 36m2. How much femcing would be needed to surround the gardens permeter if the length of the garden is three times larger than the width?"

Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

w=width
length=3w
area is their product or 3w^2
It equals 36 m^2
Therefore, 3w^2=36
w^2=12
w=2 sqrt (3) m
3w=6 sqrt (3) m
Their sum is 8 sqrt(3)
double that for the perimeter, and that is 16 sqrt (3) m

Answer by Alan3354(69443)

"A rectangulat garden has an area of 36m2. How much femcing would be needed to surround the gardens permeter if the length of the garden is three times larger than the width?"
-----------------
1 time greater = 2x as great
2 times greater = 3x as great
3 times greater = 4x as great
====================
L = W + 3W = 4W
L*W = 36
----
Sub for L
4W*W = 36
W^2 = 9
W = 3, L = 12
----
Perimeter = 2W + 2L
= 30 meters